OpenStudy (anonymous):

Compute dy/dx: x^4y^3+tany=sinx

OpenStudy (solomonzelman):

differentiate term by term

OpenStudy (solomonzelman):

what is the derivative of x^4y^3 ?

OpenStudy (solomonzelman):

(use the product rule and the chain rule for y - i.e. to multiply times dy/dx the derivative of y)

OpenStudy (anonymous):

I thought she was wanting us to use the implicit differentiation

OpenStudy (solomonzelman):

yes, we are going to

OpenStudy (solomonzelman):

I mean there is no way to go otherwise...

OpenStudy (solomonzelman):

k, do you know how to compute the derivative of x^4y^3 ?

OpenStudy (anonymous):

4x^3(3y^2)?

OpenStudy (solomonzelman):

there is a thing called a product rule for the derivative of a product of the functions f(x) times g(x). \(\large\color{black}{ \displaystyle f'(x)g(x)+f(x)g'(x) }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x^4\times y^3 ~~~_{\Huge \Longrightarrow}^{\rm becomes}~~~\color{red}{4x^3}\times y^3~+ ~x^4 \times \color{red}{3y^2 \times \frac{dy}{dx}} }\) in red are the derivatives.

OpenStudy (solomonzelman):

see how I am differentiating the product of these 2 functions?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

ok, can you differentiate the second term? (( derivative of tan(y), and don't forget that the derivative of y will be always multiplied times dy/dx ))

OpenStudy (solomonzelman):

always **be** multiplied...

OpenStudy (solomonzelman):

at first, what is the derivative of tan(x) ?

OpenStudy (anonymous):

sec^2x

OpenStudy (solomonzelman):

yes derivative of tan(x) = sec^2(x). and derivative of tan(y) = sec^2(y) \(\small\color{black}{ \displaystyle \times }\) (dy/dx)

OpenStudy (solomonzelman):

see how the derivative of y works ? (multiplying the derivative of y times dy/dx every time)

OpenStudy (solomonzelman):

so, the derivative of tan(y) is what (again) ?

OpenStudy (anonymous):

sec^2(y) haha

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

you skipped dy/dx

OpenStudy (solomonzelman):

(dy/dx) * sec^2(y)

OpenStudy (solomonzelman):

the derivative of y is same as derivative of x, BUT derivative of y is multiplied times dy/dx

OpenStudy (solomonzelman):

and what is the derivative of sin(x) ?

OpenStudy (anonymous):

cos(x)

OpenStudy (solomonzelman):

yes, so putting the terms together now...

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle ~~~~~~~~~~x^4y^3~~~~~~~~~~~~~~~~+\tan y~~~~~~~~~~=\sin x }\) \(\large\color{black}{ \displaystyle 4x^3y^3~+ ~x^4 3y^2 \color{red}{\frac{dy}{dx}}~~+\sec^2y\color{red}{\frac{dy}{dx}}~~~=\cos x }\)

OpenStudy (solomonzelman):

I put the derivatives underneath.... see ?

OpenStudy (solomonzelman):

Now, all you need to do is to isolate the dy/dx on one side.

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

can you isolate (algebraically solve for) dy/dx ?

OpenStudy (anonymous):

I'm not sure, since dy/dx isn't in all the terms. I feel like I need to thoug

OpenStudy (anonymous):

though

OpenStudy (solomonzelman):

\(\large\color{slate}{\large\color{black}{ \displaystyle 4x^3y^3~+ ~x^4 3y^2 \color{red}{\frac{dy}{dx}}~~+\sec^2y\color{red}{\frac{dy}{dx}}~~~=\cos x }}\) \(\large\color{slate}{\large\color{black}{ \displaystyle x^4 3y^2 \color{red}{\frac{dy}{dx}}~~+\sec^2y\color{red}{\frac{dy}{dx}}~~~=\cos x -4x^3y^3 }}\) \(\large\color{slate}{\large\color{black}{ \displaystyle \color{red}{\frac{dy}{dx}}\left( x^4 3y^2+\sec^2y\right)=\cos x -4x^3y^3 }}\) \(\large\color{slate}{\large\color{black}{ \displaystyle \color{red}{\frac{dy}{dx}}=\frac{\cos x -4x^3y^3}{x^4 3y^2+\sec^2y} }}\)

OpenStudy (solomonzelman):

that is all....

OpenStudy (anonymous):

Thank you

OpenStudy (solomonzelman):

yw